OpenStudy (darkigloo):

Why does ((64x)^(1/4))^2 equal (64x)^(1/2)?

jimthompson5910 (jim_thompson5910):

What are the exponents in the original expression?

OpenStudy (darkigloo):

\[((64x) ^{1/4})^{2}\]

jimthompson5910 (jim_thompson5910):

do you see the exponents?

OpenStudy (darkigloo):

yes

jimthompson5910 (jim_thompson5910):

can you list them for me?

OpenStudy (darkigloo):

1/4 and 2

jimthompson5910 (jim_thompson5910):

when you have an exponential expression raised to another exponent, you can multiply the exponents, ie use this rule \[\Large (a^b)^c = a^{b*c}\]

jimthompson5910 (jim_thompson5910):

so multiplying 1/4 and 2 gives: (1/4)*2 = (1/4)*(2/1) = (1*2)/(4*1) = 2/4 = 1/2

OpenStudy (darkigloo):

ohhh wow! thank you

jimthompson5910 (jim_thompson5910):

you're welcome